If you drop a rock off a cliff and it hits the ground 4 seconds later, how tall is that cliff?





1
Expert's answer
2020-12-17T09:13:55-0500

We can find the height of the cliff from the kinematic equation:


"y=\dfrac{1}{2}gt^2=\dfrac{1}{2}\cdot 9.8\ \dfrac{m}{s^2}\cdot (4\ s)^2=78.4\ m."

Answer:

"y=78.4\ m."